# Scale Reliant Inference

### v4

@@ 8 Discussion
In this article, we have framed the challenge of data with arbitrary scale as an estimation problem involving partially identified models. Through this formulation, we have been able to provide the first formal limits on inference with these data. Our results are consistent with the CoDA perspective: they show that care is needed as spurious scientific conclusions can be reached due to the lack of information about the system scale contained in these data. However, unlike the CoDA perspective, our results argue that a dogmatic restriction to scale invariance is overly prohibitive. Consider that, even if data is compositional (i.e., a full compositional survey with simplex valued measurements), Bayesian PIMs can be used to perform rigorous SRI. In fact, even Frequentist hypothesis testing and confidence estimation can be done rigorously when analyzing compositional data if the estimand is partially scale reliant. Overall, we argue that SRI is possible yet requires new types of statistical methods that can account for the limited information contained in the observed data.
@@
Beyond discussing limits on inference, this article also introduces the class of Bayesian PIMs for SRI. There is an interesting connection between the SSRV formulation of Bayesian PIMs and adversarial robustness. Adversarial methods are well known in the machine learning literature as a way of enhancing the robustness or stability (Yu,, 2020) of predictions. For example, within computer vision, noise is sometimes added to images in order to ensure that trained models are stable and are not sensitive to slight variation when applied out-of-sample (Santos and Papa,, 2022). Due to concerns about rigor and reproducibility in science, some authors have advocated that adversarial perturbations can be applied to ensure robustness of statistical inference (Yu,, 2020). Within SSRVs, the use of the scale model can be viewed as a type of adversarial perturbation: for each sample from the SSRV we apply a different random perturbation sampled from the scale model. One way of understanding the utility of Bayesian PIMs (and SSRVs) is that they involve a form of adversarial robustness to perturbations of scale.
@@
In this article, we have focused on Bayesian approaches to SRI as these methods, unlike Frequentist methods, remain valid even in the setting of Full SRI. To clarify the problem with Frequentist methods, consider the problem of confidence estimation. Intuitively, a level-$\alpha$ confidence set for an estimand $\theta$ should, with probability $1-\alpha$ cover the identification region ${\mbox{Id}}_{\theta}(P)$. The problem is that, for Full SRI, ${\mbox{Id}}_{\theta}(P)=\Theta$ and therefore any level-$\alpha$ confidence set must be uselessly large. Yet, for Partial SRI, useful confidence estimation (and hypothesis testing) may be possible as ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$. For example, we recently proved that useful and robust hypothesis testing is possible for a particular partial SRI problem called differential set analysis (McGovern et al.,, 2023). Overall, we believe that Frequentist SRI for Partial SRI problems is an intriguing and likely fruitful area for future research. In fact, there is already a rich literature on Frequentist confidence estimation for partially identified models can likely be adapted to the task. We refer interested readers to Imbens and Manski, (2004) and Stoye, (2009) for a thorough review of those methods.
@@
For simplicity, we have focused on compositional surveys which are a subset of data with arbitrary scale where the system composition is identifiable. Yet in many practical applications, we expect that the system composition may not be identifiable as the measurement process may involve unobserved compositional biases. For example, in the 16S rRNA microbiome surveys, the PCR amplification process can lead to an over-representation of some taxa compared to others in the sequenced pool (Silverman et al.,, 2021). If compositional bias is substantial, then it may tighten the limits on inference further restricting the types of inference that can be performed rigorously with these data. Just as Bayesian PIMs can be used to address arbitrary scale, we expect these same models may be adapted to provide robustness to compositional bias. Overall, we believe that the study of compositional bias represents a fruitful area for future research.
Acknowledgments and Funding Sources
We thank Rachel Silverman, Juan Jose Egozcue, Vera Pawlosky-Glahn, Gregory Gloor, Michael Love, and James Morton for their manuscript comments. JDS and MPN were supported in part by NIH 1R01GM148972-01. LAD acknowledges support from the the Global Probiotics Council, a Searle Scholars Award, the Hartwell Foundation, an Alfred P. Sloan Research Fellowship, the Translational Research Institute through Cooperative Agreement NNX16AO69A, the Damon Runyon Cancer Research Foundation, and NIH 1R01DK116187-01. SM would like to acknowledge the support of grants NSF IIS-1546331, NSF DMS-1418261, NSF IIS-1320357, NSF DMS-1045153, and NSF DMS-1613261.
Data Access and Code Availability
@@
All data and R scripts necessary to replicate all analyses in the manuscript and supplement are located at https://github.com/michellepistner/ScaleSimulation.

### v5

@@ 5 Discussion
In this article, we have framed the challenge of data that lacks scale information as an estimation problem involving partially identified models. Through this formulation, we have been able to provide the first formal limits on inference with these data. Our results are consistent with the CoDA perspective: they show that care is needed as spurious scientific conclusions can be reached due to the lack of information about the system scale contained in these data. However, unlike the CoDA perspective, our results argue that a dogmatic restriction to scale invariance is overly prohibitive. Consider that, even if data is compositional (i.e., a full compositional survey with simplex valued measurements), Bayesian PIMs can be used to perform SRI while satisfying Bayesian ideals. In fact, hypothesis tests and confidence intervals which satisfy Frequentist ideals can be constructed from compositional data if the estimand is partially scale reliant. Overall, we argue that SRI is possible yet requires new types of statistical methods that can account for the limited information contained in the observed data.
@@
Beyond discussing limits on inference, this article introduces the class of Bayesian PIMs for SRI. There is an interesting connection between the SSRV formulation of Bayesian PIMs and adversarial robustness. Adversarial methods are well known in the machine learning literature as a way of enhancing the robustness or stability (Yu, 2020) of predictions. For example, within computer vision, noise is sometimes added to images in order to ensure that trained models are stable and are not sensitive to slight variation when applied out-of-sample (Santos and Papa, 2022). Due to concerns about rigor and reproducibility in science, some authors have advocated that adversarial perturbations can be applied to ensure robustness of statistical inference (Yu, 2020). Within SSRVs, the use of the scale model can be viewed as a type of adversarial perturbation: for each sample from the SSRV we apply a different random perturbation sampled from the scale model. One way of understanding the utility of Bayesian PIMs (and SSRVs) is that they involve a form of adversarial robustness to perturbations of scale.
@@
In this article, we have focused on Bayesian approaches to SRI as these methods, unlike Frequentist methods, remain valid even in the setting of Full SRI. To clarify the problem with Frequentist methods, consider the problem of confidence estimation. Intuitively, a level-$\alpha$ confidence set for an estimand $\theta$ should, with probability $1-\alpha$, cover the identification region ${\mbox{Id}}_{\theta}(P)$. The problem is that, for Full SRI, ${\mbox{Id}}_{\theta}(P)=\Theta$ and therefore any level-$\alpha$ confidence set must be uselessly large. Yet, for Partial SRI, useful confidence estimation (and hypothesis testing) may be possible as ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$. For example, we recently proved that useful and robust hypothesis testing is possible for a particular partial SRI problem called differential set analysis (McGovern, Nixon and Silverman, 2023). Overall, we believe that Frequentist SRI for Partial SRI problems is an intriguing and likely fruitful area for future research. In fact, there is already a rich literature on Frequentist confidence estimation for partially identified models can likely be adapted to the task. We refer interested readers to Imbens and Manski (2004) and Stoye (2009) for a thorough review of those methods.
@@
For simplicity, we have focused on compositional surveys which are a subset of data with arbitrary scale where the system composition is identifiable. Yet in many practical applications, we expect that the system composition may not be identifiable as the measurement process may involve unobserved compositional biases. For example, in the 16S rRNA microbiome surveys, the PCR amplification process can lead to an over-representation of some taxa compared to others in the sequenced pool (Silverman et al., 2021). If compositional bias is substantial, then it may tighten the limits on inference further restricting the types of inference that can be performed rigorously with these data. Just as Bayesian PIMs can be used to address arbitrary scale, we expect these same models may be adapted to provide robustness to compositional bias. Overall, we believe that the study of compositional bias represents a fruitful area for future research.
{acks}
@@
[Data Access and Availability] All data and R scripts necessary to replicate all analyses in the manuscript and supplement are located at https://github.com/michellepistner/ScaleSimulation.
{acks}
[Acknowledgments] We thank Rachel Silverman, Juan Jose Egozcue, Vera Pawlosky-Glahn, Gregory Gloor, Michael Love, and James Morton for their manuscript comments.
{funding}
JDS and MPN were supported in part by NIH 1R01GM148972-01. LAD acknowledges support from the the Global Probiotics Council, a Searle Scholars Award, the Hartwell Foundation, an Alfred P. Sloan Research Fellowship, the Translational Research Institute through Cooperative Agreement NNX16AO69A, the Damon Runyon Cancer Research Foundation, and NIH 1R01DK116187-01. SM would like to acknowledge the support of grants NSF IIS-1546331, NSF DMS-1418261, NSF IIS-1320357, NSF DMS-1045153, and NSF DMS-1613261.
[…]